You are attempting to calculate a firm’s free cash flow to equity. You know the following information about the firm. What is the firm’s free cash flow to equity in year 2?

Year

1

2

After-tax earnings

130,000

150,000

Depreciation

30,000

25,000

Current Assets

15,000

10,000

Current Liabilities

10,000

5,000

Net PPE (this is your investment in fixed assets)

100,000

120,000

A) $145,000

B) $155,000

C) $150,000

D) $140,000

E) $135,000
Answer #1

increase in working capital = working capital year 2 - working capital year 1 = (Current asset year 2 - current liability year 2) - (Current asset year 1 - current liability year 1)= (15000-10000)-(10000-5000)=0

increase in net ppe = (net ppe year 2 - net ppe year 1) = (120000-100000)=20000

Free cash flow = after tax earning + depreciation - increase in working capital - increase in net ppe = 150000+25000-0-20000=$155000 (opt b)